DIAGNOSTICS AND SCORE STATISTICS IN REGRESSION

P. C. WANG

Graduate Institute of Statistics, National Central University, Chung-Li, Taiwan 32054, Republic of China

(Received June 27, 1989; revised June 21, 1990)

Abstract. In a generalized linear model, we have a linear predictor. We extend to a nonlinear one and propose a unified method to establish diagnostic procedures for such models with nonlinear links. Applications of the procedures to various useful models are given with examples.

Key words and phrases: Generalized linear model, nonlinear, diagnostic plot, case influence.